OpenStudy (anonymous):

How would I solve this:? A=2P-3R for R

OpenStudy (anonymous):

subtract \(2P\) then divide by \(-3\)

OpenStudy (anonymous):

how would the whole problem be worked out?

OpenStudy (anonymous):

\[A=2P-3R\] \[A-2P=2P-2P-3R\] \[A-2P=-3R\] \[\frac{A-2P}{-3}=\frac{-3R}{-3}\] \[\frac{A-2P}{-3}=R\] or maybe write \[\frac{2P-A}{3}=R\]